18x^2+12xy-3rx-2ry
Found 2 solutions by fractalier, stanbon:
Answer by fractalier(2101)

Factor by grouping...
18x^2 + 12xy - 3rx - 2ry =
(18x^2 + 12xy) - (3rx + 2ry) =
6x(3x + 2y) - r(3x + 2y) =
(6x - r)(3x + 2y)

Answer by stanbon(48579)

18x^2+12xy-3rx-2ry
=6x(3x+2y)-r(3x+y)
=(3x+y)(6x-r)
cheers,
stan H.